OpenStudy (anonymous):

What is the equation of the line that passes through the point (3, 2) and has a slope of 4? y = 4x - 1 y = 4x - 5 y = 4x - 10 y = 4x + 1

OpenStudy (anonymous):

here we go again right?

OpenStudy (anonymous):

Yes. I know>.< I forgot AGAIN.

OpenStudy (anonymous):

once more it is the point- slope formula \[y-y_1=m(x-x_1)\] make a direct substitution with the numbers you have

OpenStudy (anonymous):

that is \(x_1=3,y_1=2,m=4\) let me know what you get

OpenStudy (anonymous):

write it on your keds before the test

OpenStudy (anonymous):

keds?

OpenStudy (anonymous):

general form of the line is y=mx+c, m is the slope and c is the y-intercept since you are given that m=4,apply it to the equation gives you y=4x+c since (3,2) is on the line,substitute into the equation gives you, 3=4(2)+c .'. c=-5 .'. y=4x-5